f is a linear function, f (x) = mx + c. a. Find f (x) if it is equal to its inverse: f ‐ ¹ (x) = f (x). b. Find f (x), if f (f (x)) = f (x)

This problem has got two solutions.


1
Expert's answer
2021-02-04T07:34:14-0500

a. If "f" maps "x" to "y," then "f^{-1}" maps "y" to "x."This gives rise to the cancellation formulas:

"f^{-1}(f(x))=x," ​for every "x" in the domain of"f(x),"

"f(f^{-1}(x))=x," ​for every "x" in the domain of"f^{-1}(x)."

Given "f^{-1}(f(x))=f(x)." Then "f(x)=x."


b. Given "f(f(x))=f(x)." Then


"m(mx+b)+b=mx+b"

"m^2x+(m+1)b=mx+b"

"m^2=m""m+1=1\ or\ b=0"



Then "f(x)=b" or "f(x)=x."